### v1

@@ 4 Transformer model
The one-layer transformer model is parameterized by the following matrices and vectors:
@@
  • Let $Z\in\mathbb{R}^{d\times\omega}$. This is the “embedding matrix.” Its $\gamma$th column corresponds to the embedding of token $\gamma$ in $\mathbb{R}^{d}$. We call $d$ the “embedding dimension.”
@@
  • Let $U\in\mathbb{R}^{d\times T}$. This is the “position matrix.” While we think of it as having only $T$ columns, it can actually be extended to arbitrarily many columns by appending new columns whenever an input with length greater than $T$ is encountered.
@@
  • Let $W_{1,r},W_{2,r}\in\mathbb{R}^{d\times d_{r}},W_{3,r}\in\mathbb{R}^{d\times d_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}\in\mathbb{R}^{m_{0}d_{0}\times d}$. These are the “self-attention matrices.” We call $m_{0}$ the “number of heads.”
@@
  • Let $W\in\mathbb{R}^{d\times m},b\in\mathbb{R}^{m},V\in\mathbb{R}^{m\times\omega}$. These are the “hidden matrix,” “bias vector,” and “output matrix” respectively. We call $m$ the “number of neurons.”
@@
Additionally, let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at some point $\eta\in\mathbb{R}$ and not a polynomial there. This is the “activation function.” For example, $\psi$ could be a sigmoid, such as the logistic function, tanh, or arctan, or it could be a smoothed rectified linear unit, such as GELU.
Let $\theta$ denote the concatenation of all the model parameters. Then, for all $\tau\in\mathbb{N}$ and $\alpha\in[\omega]^{\tau}$, define
@@
$\begin{split}&h_{1,\theta}(\alpha)=Z[:,\alpha]+U[:,1:\tau]\in\mathbb{R}^{d\times\tau}\\ &h_{2,\theta}(\alpha)=W_{0}^{\top}\text{vec}\left(\left[W_{3,r}^{\top}h_{1,\theta}(\alpha)\phi\left(h_{1,\theta}(\alpha)^{\top}W_{1,r}W_{2,r}^{\top}h_{1,\theta}(\alpha)[:,\tau]\right)\right]_{r\in[m_{0}]}\right)\in\mathbb{R}^{d}\\ &h_{\theta}(\alpha)=V^{\top}\psi\left(W^{\top}h_{2,\theta}(\alpha)+b\right)\in\mathbb{R}^{\omega}.\end{split}$ (2)
@@
These three functions are the “embedding,” “self-attention,” and “FNN” sub-layers respectively of a single transformer layer. Note that we do not include the skip-connection $+h_{1,\theta}(\alpha)[\tau]$ in the self-attention sub-layer. This was omitted for simplicity, but its inclusion does not affect Theorems 4 and 5, which follow by almost the same proof.
5 The rank results
First, we need to derive the generic rank of $\psi(ab^{\top})$. We will do this for polynomial $\psi$ and then extend to $\psi$ which are real analytic at zero.
Theorem 2.
@@ 5 The rank results
To see that $M$ is open, notice that it is the pre-image of $(-\rho,\rho)^{m\times n}$ under the continuous map $a\mapsto ab^{\top}$. Now, let $(c_{k})$ be the coefficients of the Taylor expansion at zero of $\psi$. Define $K=\{k\in\mathbb{N}\cup\{0\}\mid c_{k}\neq 0\}$. Define $s=\min\{m,n\}$ and let $k_{1},\ldots,k_{s}$ be the first $s$ integers in $K$. Define $K_{s}=\{k_{1},\ldots,k_{s}\}$. Let $I\subset[m]$ and $J\subset[n]$ such that $|I|=|J|=s$. Define $p:M\to\mathbb{R}:a\mapsto\det_{I,J}(\sum_{k\in K_{s}}c_{k}(ab^{\top})^{(k)})$ and $q:M\to R:a\mapsto\det_{I,J}(\psi(ab^{\top}))$. Note from the proof of Theorem 2 that the monomials of $p$ all have degree $k_{1}+\cdots+k_{s}$ whereas if $S\in\binom{K}{s}$ has any integers not in $K_{s}$, then its corresponding monomials will have degree $>k_{1}+\cdots+k_{s}$. Thus, $p$ is the truncation of the Taylor expansion at zero of $q$ to precisely the monomials of degree $k_{1}+\cdots+k_{s}$. Since, by Theorem 2, $p$ is not identically zero, it has at least one nonzero coefficient, and so the Taylor expansion at zero of $q$ has at least one nonzero coefficient. Thus, $q$ is not identically zero, completing the proof. ∎
6 The memory capacity results
With Theorem 3 in hand, we are ready to prove the following fact about self-attention with embedding dimension 1.
Theorem 4.
Let $\omega,\tau\in\mathbb{N}$. Let $\phi$ be the softmax function. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\omega}\times\mathbb{R}^{\tau}\times\cup_{t=1}^{\tau}[\omega]^{t}\to\mathbb{R}$
$(z,u,\alpha)\mapsto\langle x,\phi(xx_{t})\rangle$
$\text{where }t=\mathrm{len}(\alpha),x=z[\alpha]+u[1:t].$
@@ 6 The memory capacity results
First, let $A$ denote $\cup_{t=1}^{\tau}[\omega]^{t}$. Then, note that, for all $z\in\mathbb{R}^{\omega},u\in\mathbb{R}^{\tau},\sigma\in[\tau],\alpha\in[\omega]^{\sigma}$,
$f(z,u,\alpha)=\frac{\sum_{t=1}^{\sigma}(z_{\alpha_{t}}+u_{t})\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\sigma}}+u_{\sigma}))}{\sum_{t=1}^{\sigma}\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\sigma}}+u_{\sigma}))}.$
To show that $\alpha$ and $\beta$ have the same length, set $z=0$ and $u=(1,\ldots,\tau)$, and define the polynomials
$p_{\gamma}=\sum_{t=1}^{\mathrm{len}(\gamma)}tx^{t\mathrm{len}(\gamma)}\text{ and }q_{\gamma}=\sum_{t=1}^{\mathrm{len}(\gamma)}x^{t\mathrm{len}(\gamma)}\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}.$
Note that the monomial $x^{\mathrm{len}(\alpha)^{2}+\mathrm{len}(\beta)^{2}}$ has coefficient $\mathrm{len}(\alpha)-\mathrm{len}(\beta)$ in $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}$. Also, $f(z,u,\gamma)=p_{\gamma}(\exp(1))/q_{\gamma}(\exp(1))\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}$. So, since $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)\equiv 0$, $(p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha})(\exp(1))=0$. But, $\exp(1)$ is transcendental, so $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}\equiv 0$. Thus, $\mathrm{len}(\alpha)=\mathrm{len}(\beta)$.
Let $\sigma=\mathrm{len}(\alpha)=\mathrm{len}(\beta)$. Let $s\in[\sigma]$. To show that $\alpha_{s}=\beta_{s}$, set $z=(1,\ldots,\omega)$ and $u=\omega e_{s}$, and define the polynomials
$p_{\gamma}=\sum_{t=1}^{\sigma}(\gamma_{t}+\omega\delta_{t,s})x^{(\gamma_{t}+\omega\delta_{t,s})(\gamma_{\sigma}+\omega\delta_{\sigma,s})}\text{ and }q_{\gamma}=\sum_{t=1}^{\sigma}x^{(\gamma_{t}+\omega\delta_{t,s})(\gamma_{\sigma}+\omega\delta_{\sigma,s})}\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}.$
Note that the monomial $x^{(\alpha_{s}+\omega)(\alpha_{\sigma}+\omega\delta_{\sigma,s})+(\beta_{s}+\omega)(\beta_{\sigma}+\omega\delta_{\sigma,s})}$ has coefficient $\alpha_{s}-\beta_{s}$ in $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}$. Also, $f(z,u,\gamma)=p_{\gamma}(\exp(1))/q_{\gamma}(\exp(1))\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}$. So, again, $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}\equiv 0$. Thus, $\alpha_{s}=\beta_{s}$. Since this was for arbitrary $s\in[\sigma]$, we get that $\alpha=\beta$, proving the theorem. ∎
It is natural to wonder what would happen if we replaced $\phi$ with the function $x\mapsto a^{x}/\sum_{\gamma=1}^{\omega}a^{x_{\gamma}}$ for some $a>1$ which is not transcendental. In this case, since $a^{b}=\exp(b\log(a))$, we would just need to scale our example $(z,u)$’s by $\sqrt{\log(a)}$ to get the same memory capacity result.
Combining Theorem 3 and Theorem 4 leads to a memory capacity lower bound for one-layer decoder-only transformers.
Theorem 5.
Let $d,m_{0},d_{0},m,\omega,T\in\mathbb{N}$ and $d_{r}\in\mathbb{N}\leavevmode\nobreak\ \forall r\in[m_{0}]$. Let $\phi$ be the softmax function and let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Let $h_{\theta}$ be the one-layer transformer model defined in Equation (LABEL:eq:model). Then the memory capacity of $\{\phi\circ h_{\theta}\mid\theta\}$ is greater than or equal to $m$.
Proof.
We will reduce to the setting where $d=m_{0}=d_{0}=d_{1}=1$. Given $z\in\mathbb{R}^{\omega},u\in\mathbb{R}^{T},w\in\mathbb{R}^{m}$, define $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d},U=\mathbbm{1}_{d}u^{\top}/\sqrt{d},W_{1,r}=W_{2,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{r}}^{\top}/\sqrt{dd_{r}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{3,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{0}}^{\top}/\sqrt{dd_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}=\mathbbm{1}_{m_{0}d_{0}}\mathbbm{1}_{d}^{\top}/(m_{0}\sqrt{dd_{0}}),W=\mathbbm{1}_{d}w^{\top}/\sqrt{d},b=\eta\mathbbm{1}_{m}$. Then, for all $\alpha\in\cup_{t=1}^{T}[\omega]^{t}$, $W^{\top}h_{2,\theta}(\alpha)=wf(z,u,\alpha)$, where we use the definition of $f$ from Theorem 4.
@@
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{T}[\omega]^{t}$ be distinct. Invoking Theorem 4, choose $(z,u)$ such that $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Invoking Theorem 3, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y\in\mathbb{R}^{\omega\times n}$. Then, there exists $V\in\mathbb{R}^{m\times\omega}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$. Since $\phi(\mathbb{R}^{\omega})=\mathrm{ri}\Delta^{\omega-1}$, the result follows. ∎
In the proof, token $\gamma\in[\omega]$ is embedded as $z_{\gamma}\mathbbm{1}_{d}/\sqrt{d}$ and the $t$th position vector is $u_{t}\mathbbm{1}_{d}/\sqrt{d}$. Thus, the inputs to the self-attention sub-layer are sequences of vectors coming from a one-dimensional subspace, so it makes sense for the self-attention matrices to be constant matrices. Finally, the $k$th hidden neuron of the FNN sub-layer has weight vector $w_{k}\mathbbm{1}_{d}/\sqrt{d}$.
@@
Theorem 5 also suggests an algorithm for finding an interpolating solution. First, note that every minor of $\psi(W^{\top}[h_{2,\theta}(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$ is real analytic on an open neighborhood of zero. Thus, it works to (1) sample $\theta$ from any continuous probability distribution, (2) choose $\epsilon>0$ sufficiently small so that the entries of $\epsilon W^{\top}[h_{2,\theta}(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top}$ are each in the interval of convergence of $\psi$, and (3) solve for $V$ in $Y=V^{\top}\psi(W^{\top}[h_{2,\theta}(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$.
@@
To compare the lower bound of Theorem 5 to the upper bound of Lemma 1, let $d_{r}=d_{1}\leavevmode\nobreak\ \forall r\in[m_{0}]$ for simplicity. Then, the total number of parameters is $\omega m+m(d+1)+2m_{0}(d_{0}+d_{1})d+(\omega+T)d$. Thus, the ratio between the upper and lower bounds is

### v2

@@ 5 Transformer model
The one-layer multi-head decoder-only transformer model is parameterized by the following matrices and vectors:
@@
  • Let $Z\in\mathbb{R}^{d\times\omega}$. This is the embedding matrix. Its $\gamma$th column corresponds to the embedding of token $\gamma$ in $\mathbb{R}^{d}$. We call $d$ the embedding dimension.
@@
  • Let $U\in\mathbb{R}^{d\times\mathbb{N}}$. This is the position matrix. While we assume it has infinite columns, it can actually start with finite columns and be extended to arbitrarily many columns by appending new columns whenever a longer input is encountered.
@@
  • Let $W_{1,r},W_{2,r}\in\mathbb{R}^{d\times d_{r}},W_{3,r}\in\mathbb{R}^{d\times d_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}\in\mathbb{R}^{m_{0}d_{0}\times d}$. These are the self-attention matrices. We call $m_{0}$ the number of heads.
@@
  • Let $W\in\mathbb{R}^{d\times m},b\in\mathbb{R}^{m},V\in\mathbb{R}^{m\times\omega}$. These are the hidden matrix, bias vector, and output matrix respectively. We call $m$ the number of neurons.
@@
Additionally, let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at a point and not a polynomial there. This is the activation function. For example, $\psi$ could be a sigmoid, such as the logistic function, tanh, or arctan, or it could be a smoothed rectified linear unit, such as GELU.
Let $\theta$ denote the concatenation of all the model parameters. Define
@@
$\begin{split}&h_{1,\theta}:\cup_{t=1}^{\infty}[\omega]^{t}\to\cup_{t=1}^{\infty}\mathbb{R}^{d\times t}:\alpha\mapsto Z[:,\alpha]+U[:,1:|\alpha|],\\ &h_{2,\theta}:\cup_{t=1}^{\infty}\mathbb{R}^{d\times t}\to\mathbb{R}^{d}:X\mapsto W_{0}^{\top}\text{vec}\left(\left[W_{3,r}^{\top}X\phi\left(X^{\top}W_{1,r}W_{2,r}^{\top}X[:,-1]\right)\right]_{r\in[m_{0}]}\right),\\ &h_{3,\theta}:\mathbb{R}^{d}\to\mathbb{R}^{\omega}:x\mapsto V^{\top}\phi\left(W^{\top}x+b\right),\\ \text{and }&h_{\theta}\coloneqq\phi\circ h_{3,\theta}\circ h_{2,\theta}\circ h_{1,\theta}\end{split}$ (6)
@@
where $-1$ denotes the last index and $\phi$ is the softmax function. We think of the value $h_{\theta}(\leavevmode\nobreak\ )$ as an additional $\omega-1$ model parameters. The first three functions are the embedding, self-attention, and FNN sub-layers respectively of the one-layer transformer $h_{\theta}$. Note that we do not include the skip-connection $+X[:,-1]$ in the self-attention sub-layer. This was omitted for simplicity, but its inclusion does not affect Theorems 6.1 and 6.5, which follow by almost the same proof.
6 Transformer next-token prediction capacity
In order to lower bound the next-token prediction capacity of the transformer model, we first reduce to the special case of $d=d_{1}=d_{0}=m_{0}=1$ and $W_{1,1}=W_{2,1}=W_{3,1}=W_{0}=1$. Thus, the model becomes the following composition of mappings:
[…]
where $f(z,u,\alpha)=\langle x,\phi(xx_{|\alpha|})\rangle$ with $x=z[\alpha]+u[1:|\alpha|]$. Assume $\psi:\mathbb{R}\to\mathbb{R}$ is real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Assume $m\geq n$. We will show that this simplified model can interpolate arbitrary data sets of size $n$. Then we will use this to show that the full model can interpolate arbitrary data sets of size $n$.
To show the former, let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\omega]^{t}$ and $Y^{\prime}\in(\mathrm{ri}\Delta^{\omega-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\omega\times n}$ such that $\phi(Y)=Y^{\prime}$ with $\phi$ applied column-wise. We will construct an interpolating model. First, set $b=\eta\mathbbm{1}_{m}$. Second, define $T=\max_{i\in[n]}|\alpha^{i}|$ and sample $(z,u)\in\mathbb{R}^{\omega}\times\mathbb{R}^{T}$ from any continuous probability distribution. Let $\hat{x}_{1},\ldots,\hat{x}_{n}$ be the corresponding outputs of the first mapping. Then the matrix of outputs of the composition of mappings is $\phi(V^{\top}\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top}))$. In Theorem 6.1, we prove that the $\hat{x}_{i}$ are nonzero and distinct. In Theorem 6.4, we prove that this implies $\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has full rank for generic $w$ in an open set around zero. So, as our third step in the construction, sample $w\in\mathbb{R}^{m}$ from any continuous probability distribution that has support in the open set. Then, $\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ has rank $n$ so, as our fourth and final step in the construction, solve for $V$ in $Y=V^{\top}\psi(w[\hat{x}_{i}]_{i\in[n]}^{\top}+b\mathbbm{1}_{n}^{\top})$ as a system of linear equations.
Thus, the simplified model is able to interpolate arbitrary data sets as long as $m\geq n$. To see how to reduce the full model with $d,m_{0},d_{0}\in\mathbb{N}$ and $d_{r}\in\mathbb{N}\leavevmode\nobreak\ \forall r\in[m_{0}]$ to the simplified model, set $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d},U=\mathbbm{1}_{d}u^{\top}/\sqrt{d},W_{1,r}=W_{2,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{r}}^{\top}/\sqrt{dd_{r}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{3,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{0}}^{\top}/\sqrt{dd_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}=\mathbbm{1}_{m_{0}d_{0}}\mathbbm{1}_{d}^{\top}/(m_{0}\sqrt{dd_{0}}),W=\mathbbm{1}_{d}w^{\top}/\sqrt{d}$. In other words, token $\gamma\in[\omega]$ is embedded as $z_{\gamma}\mathbbm{1}_{d}/\sqrt{d}$ and the $t$th position vector is $u_{t}\mathbbm{1}_{d}/\sqrt{d}$. So, the inputs to the self-attention sub-layer are sequences of vectors coming from a one-dimensional subspace, and so it makes sense for the self-attention matrices to be constant matrices. Finally, the $k$th hidden neuron of the FNN sub-layer has weight vector $w_{k}\mathbbm{1}_{d}/\sqrt{d}$. By reducing to the simplified model in this way, and applying Theorems 6.1 and 6.4, we are able to lower bound the next-token prediction capacity of the one-layer multi-head decoder-only transformer model by $m$ in Theorem 6.5.
To summarize the proof sketch, each sub-layer has a particular role in the proof and these roles are:
  1. 1.
    the embedding and self-attention sub-layers map distinct sequences to distinct vectors,
  2. 2.
    the hidden FNN layer maps these vectors to a larger space where they are linearly independent,
  3. 3.
    the outer FNN layer is solved as a system of linear equations.
6.1 Injectivity of self-attention
To prove that the embedding and self-attention layers can take on the first role in the proof sketch, we have the following theorem.
Theorem 6.1.
Let $\omega,T\in\mathbb{N}$. Let $\phi$ be the softmax function. Define
$f:\leavevmode\nobreak$ $\mathbb{R}^{\omega}\times\mathbb{R}^{T}\times\cup_{t=1}^{T}[\omega]^{t}\to\mathbb{R}$
$(z,u,\alpha)\mapsto\langle x,\phi(xx_{|\alpha|})\rangle$
$\text{where }x=z[\alpha]+u[1:|\alpha|].$
Proof.
First, let $A$ denote $\cup_{t=1}^{T}[\omega]^{t}$. Then, note that, for all $z\in\mathbb{R}^{\omega},u\in\mathbb{R}^{T},\tau\in[T],\alpha\in[\omega]^{\tau}$,
$f(z,u,\alpha)=\frac{\sum_{t=1}^{\tau}(z_{\alpha_{t}}+u_{t})\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\tau}}+u_{\tau}))}{\sum_{t=1}^{\tau}\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\tau}}+u_{\tau}))}.$
To show that $\alpha$ and $\beta$ have the same length, set $z=0$ and $u=(1,\ldots,T)$, and define the polynomials
$p_{\gamma}=\sum_{t=1}^{|\gamma|}tx^{t|\gamma|}\text{ and }q_{\gamma}=\sum_{t=1}^{|\gamma|}x^{t|\gamma|}\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}.$
Note that the monomial $x^{|\alpha|^{2}+|\beta|^{2}}$ has coefficient $|\alpha|-|\beta|$ in $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}$. Also, for all $\gamma\in\{\alpha,\beta\}$, $f(z,u,\gamma)=p_{\gamma}(\exp(1))/q_{\gamma}(\exp(1))$. So, since $f(\cdot,\cdot,\alpha)-f(\cdot,\cdot,\beta)\equiv 0$, $(p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha})(\exp(1))=0$. But, $\exp(1)$ is transcendental, so $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}\equiv 0$. Thus, $|\alpha|=|\beta|$.
Let $\tau=|\alpha|=|\beta|$. Let $s\in[\tau]$. To show that $\alpha_{s}=\beta_{s}$, set $z=(1,\ldots,\omega)$ and $u=\omega e_{s}$, and define the polynomials
$p_{\gamma}=\sum_{t=1}^{\tau}(\gamma_{t}+\omega\delta_{t,s})x^{(\gamma_{t}+\omega\delta_{t,s})(\gamma_{\tau}+\omega\delta_{\tau,s})}\text{ and}q_{\gamma}=\sum_{t=1}^{\tau}x^{(\gamma_{t}+\omega\delta_{t,s})(\gamma_{\tau}+\omega\delta_{\tau,s})}\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}.$
Note that the monomial $x^{(\alpha_{s}+\omega)(\alpha_{\tau}+\omega\delta_{\tau,s})+(\beta_{s}+\omega)(\beta_{\tau}+\omega\delta_{\tau,s})}$ has coefficient $\alpha_{s}-\beta_{s}$ in $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}$. Also, $f(z,u,\gamma)=p_{\gamma}(\exp(1))/q_{\gamma}(\exp(1))\leavevmode\nobreak\ \forall\gamma\in\{\alpha,\beta\}$. So, again, $p_{\alpha}q_{\beta}-p_{\beta}q_{\alpha}\equiv 0$. Thus, $\alpha_{s}=\beta_{s}$. Since this was for arbitrary $s\in[\tau]$, we get that $\alpha=\beta$, proving the theorem. ∎
After assuming $f(\cdot,\cdot,\alpha)\equiv f(\cdot,\cdot,\beta)$, the main idea of the rest of the proof was that
$\frac{\sum_{t=1}^{\tau}(z_{\alpha_{t}}+u_{t})\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\tau}}+u_{\tau}))}{\sum_{t=1}^{\tau}\exp((z_{\alpha_{t}}+u_{t})(z_{\alpha_{\tau}}+u_{\tau}))}=\frac{\sum_{t=1}^{\tau}(z_{\beta_{t}}+u_{t})\exp((z_{\beta_{t}}+u_{t})(z_{\beta_{\tau}}+u_{\tau}))}{\sum_{t=1}^{\tau}\exp((z_{\beta_{t}}+u_{t})(z_{\beta_{\tau}}+u_{\tau}))}$
for all $(z,u)$, so when we multiply denominators and move everything to one side, we get that $\exp(1)$ is the root of a polynomial. Since $\exp(1)$ is transcendental, the polynomial is the zero polynomial, so each monomial has coefficient equal to zero. Thus, the idea is to choose $(z,u)$ such that the coefficient of a particular monomial being zero is precisely what we are trying to show. Moreover, we can do this for as many $(z,u)$’s as we want. First, we set $z=0$ and $u=(1,\ldots,T)$, targeting the lengths of $\alpha$ and $\beta$. Here, it makes sense that only the positional embedding is needed. Then, we set $z=(1,\ldots,\omega)$ and $u=\omega e_{s}$, using the positional embedding to target position $s$. These choices of $(z,u)$ turn out to be sufficient to show that $\alpha=\beta$.
Remark 6.2.
It is quite interesting that we needed the transcendence of $\exp(1)$ for the proof and it is natural to wonder what would happen if we replaced $\phi$ with the function $x\mapsto a^{x}/\sum_{\gamma=1}^{\omega}a^{x_{\gamma}}$ for some $a>1$ which is not transcendental. In this case, since $a^{b}=\exp(b\log(a))$, we would just need to scale our example $(z,u)$’s by $\sqrt{\log(a)}$ to get the same next-token prediction capacity result. So, the important aspect of self-attention that leads to its injectivity is that $\phi$ has the form $x\mapsto a^{x}/\sum_{\gamma=1}^{\omega}a^{x_{\gamma}}$, not that the base happens to be $\exp(1)$.
It is also natural to wonder what is the simplest alternative to self-attention that retains injectivity. We will explore this question later in Section 8, but for now we continue with the next-token prediction capacity of the transformer model.
6.2 Rank results
To prove that the hidden FNN layer can take on the second role in the proof sketch, we have the following two theorems. In the first theorem, we derive the generic rank of $\psi(ab^{\top})$ for polynomial $\psi$. In the second theorem, we extend to $\psi$ which are real analytic at zero and not a polynomial there.
Theorem 6.3.
@@ 6.2 Rank results
To see that $M$ is open, notice that it is the pre-image of $(-\rho,\rho)^{m\times n}$ under the continuous map $a\mapsto ab^{\top}$. Now, let $(c_{k})$ be the coefficients of the Taylor expansion at zero of $\psi$. Define $K=\{k\in\mathbb{N}\cup\{0\}\mid c_{k}\neq 0\}$. Define $s=\min\{m,n\}$ and let $k_{1},\ldots,k_{s}$ be the first $s$ integers in $K$. Define $K_{s}=\{k_{1},\ldots,k_{s}\}$. Let $I\subset[m]$ and $J\subset[n]$ such that $|I|=|J|=s$. Define $p:M\to\mathbb{R}:a\mapsto\det_{I,J}(\sum_{k\in K_{s}}c_{k}(ab^{\top})^{(k)})$ and $q:M\to R:a\mapsto\det_{I,J}(\psi(ab^{\top}))$. Note from the proof of Theorem 6.3 that the monomials of $p$ all have degree $k_{1}+\cdots+k_{s}$ whereas if $S\in\binom{K}{s}$ has any integers not in $K_{s}$, then its corresponding monomials will have degree $>k_{1}+\cdots+k_{s}$. Thus, $p$ is the truncation of the Taylor expansion at zero of $q$ to precisely the monomials of degree $k_{1}+\cdots+k_{s}$. Since, by Theorem 6.3, $p$ is not identically zero, it has at least one nonzero coefficient, and so the Taylor expansion at zero of $q$ has at least one nonzero coefficient. Thus, $q$ is not identically zero, completing the proof. ∎
6.3 Interpolation from the FNN
Combining Theorem 6.1 and Theorem 6.4 leads to a next-token prediction capacity lower bound for one-layer multi-head decoder-only transformers.
Theorem 6.5.
Let $d,m_{0},d_{0},m,\omega\in\mathbb{N}$ and $d_{r}\in\mathbb{N}\leavevmode\nobreak\ \forall r\in[m_{0}]$. Let $\psi:\mathbb{R}\to\mathbb{R}$ be real analytic at some $\eta\in\mathbb{R}$ and not a polynomial there. Let $h_{\theta}$ be the one-layer multi-head decoder-only transformer model defined in Equation (LABEL:eq:model). Then the next-token prediction capacity of $\{h_{\theta}\mid\theta\}$ is greater than or equal to $m$.
@@ 6.3 Interpolation from the FNN
Set $n=m$. Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\omega]^{t}$ be distinct and define $T=\max_{i\in[n]}|\alpha^{i}|$. Invoking Theorem 6.1, choose $(z,u)$ such that $[f(z,u,\alpha^{i})]_{i\in[n]}$ has entries which are nonzero and distinct. Invoking Theorem 6.4, choose $w$ such that $\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$ has rank $n$. Let $Y^{\prime}\in(\mathrm{ri}\Delta^{\omega-1})^{n}$. Then there exists $Y\in\mathbb{R}^{\omega\times n}$ such that $\phi(Y)=Y^{\prime}$ where $\phi$ is the column-wise softmax function. Furthermore, there exists $V\in\mathbb{R}^{m\times\omega}$ such that $Y=V^{\top}\psi(w[f(z,u,\alpha^{i})]_{i\in[n]}^{\top}+\eta\mathbbm{1}_{m}\mathbbm{1}_{n}^{\top})$. So, by defining $Z=\mathbbm{1}_{d}z^{\top}/\sqrt{d},U=\mathbbm{1}_{d}u^{\top}/\sqrt{d},W_{1,r}=W_{2,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{r}}^{\top}/\sqrt{dd_{r}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{3,r}=\mathbbm{1}_{d}\mathbbm{1}_{d_{0}}^{\top}/\sqrt{dd_{0}}\leavevmode\nobreak\ \forall r\in[m_{0}],W_{0}=\mathbbm{1}_{m_{0}d_{0}}\mathbbm{1}_{d}^{\top}/(m_{0}\sqrt{dd_{0}}),W=\mathbbm{1}_{d}w^{\top}/\sqrt{d},b=\eta\mathbbm{1}_{m}$, we get that $Y^{\prime}=[h_{\theta}(\alpha^{i})]_{i\in[n]}$, completing the proof. ∎
Theorem 6.5 shows how the one-layer transformer model can interpolate $m$ input-output pairs through the following mechanism: (1) the embedding and self-attention sub-layers map distinct sequences to distinct vectors, (2) the hidden FNN layer maps these vectors to a larger space where they are linearly independent, (3) the outer FNN layer is solved as a system of linear equations.
@@
Theorem 6.5 also suggests an algorithm for finding an interpolating solution. First, note that every minor of $\psi(W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$ is real analytic on an open neighborhood of zero. Thus, it works to (1) sample $\theta$ from any continuous probability distribution, (2) choose $\epsilon>0$ sufficiently small so that the entries of $\epsilon W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top}$ are each in the interval of convergence of $\psi$, and (3) solve for $V$ in $Y=V^{\top}\psi(W^{\top}[(h_{2,\theta}\circ h_{1,\theta})(\alpha^{i})]_{i\in[n]}+b\mathbbm{1}_{n}^{\top})$.
@@
To compare the lower bound of Theorem 6.5 to the upper bound of Lemma 4.2, we have to assume a maximum context length $T$. Also, let $d_{r}=d_{1}\leavevmode\nobreak\ \forall r\in[m_{0}]$ for simplicity. Then, the total number of parameters is $\omega m+m(d+1)+2m_{0}(d_{0}+d_{1})d+(\omega+T)d$. Thus, the ratio between the upper and lower bounds is